# Marginal treatment effects in the absence of instrumental variables

### v2

@@ 4 Estimation
The kernel-weighted pairwise difference estimator has the advantage of having a closed-form expression, so we need not solve any formidable optimization problems. However, it faces the challenging problem of bandwidth selection, similar to most semiparametric estimation methods. Alternatively, we can consider imposing a parametric specification on the unobservable heterogeneity of the MTE such that $E\left[\left.U_{d}\right|V=v\right]=E\left[\left.U_{d}\right|V=v;\theta_{d}\right]$ for finite dimensional $\theta_{d}$, e.g., the polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}v^{j}$ or the normal polynomial specification that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}\Phi^{-j}\left(v\right)$. In the latter, setting $J=1$ will match Heckman’s normal sample selection model. Under the parametric restriction, the second step becomes a global regression $E\left[Y\left|X,D=d\right.\right]=X^{\prime}\beta_{d}+g_{d}\left(P;\theta_{d}\right)$ with parameterized selection bias correction term $g_{d}\left(p;\theta_{d}\right)$ for each $d$; therefore, the tuning of $h_{2}$ and $h_{3}$ is circumvented. Another advantage of a parametrically specified second step is the lower computational burden compared with that of the pairwise difference estimator defined by double summation which requires a squared amount of calculation (Pan and Xie, 2023).
@@
$E\left[Y\left|X\right.\right]$ $=$ $E\left[Y_{0}+D\left(Y_{1}-Y_{0}\right)\left|X\right.\right]$ (13)
$=$ $E\left[Y_{0}\left|X\right.\right]+E\left[\left.Y_{1}-Y_{0}\right|X,D=1\right]\Pr\left(D=1\left|X\right.\right)$
@@
Note that since
$q^{\left(1\right)}\left(p\right)\equiv\frac{dq\left(p\right)}{dp}=E\left[U_{1}-U_{0}\left|V=p\right.\right],$
@@
$\left(\begin{array}[]{c}\hat{\beta}_{0}\\ \hat{\delta}\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)^{\prime}\right]^{-1}$
@@
where $\hat{\delta}$ is an estimator for $\beta_{1}-\beta_{0}$ according to (13). Given $\hat{\beta}_{0}$ and $\hat{\delta}$, we apply the local linear method as well, yielding
$\left(\begin{array}[]{c}\hat{r}\left(p\right)\\ \hat{q}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)^{\prime}\right]^{-1}$
$\cdot\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\hat{\beta}_{0}-\hat{P}_{i}X_{i}^{\prime}\hat{\delta}\right)\right],$
where $\hat{r}\left(p\right)$ is an estimator for $\alpha_{0}+q\left(p\right)$. Then, we construct the instrument-free LIV estimators for the MTE and other treatment parameters as
$\hat{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\hat{\delta}+\hat{q}^{\left(1\right)}\left(v\right),$
$\hat{\Delta}^{\text{ATE}}\left(x\right)$ $=$ $x^{\prime}\hat{\delta}+\hat{r}\left(1\right)-\hat{r}\left(0\right),$
$\hat{\Delta}^{\text{TT}}\left(x\right)$ $=$ $x^{\prime}\hat{\delta}+\frac{\hat{r}\left(\hat{\pi}\left(x\right)\right)-\hat{r}\left(0\right)}{\hat{\pi}\left(x\right)},$
$\hat{\Delta}^{\text{TUT}}\left(x\right)$ $=$ $x^{\prime}\hat{\delta}+\frac{\hat{r}\left(1\right)-\hat{r}\left(\hat{\pi}\left(x\right)\right)}{1-\hat{\pi}\left(x\right)},$
$\hat{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $x^{\prime}\hat{\delta}+\frac{\hat{r}\left(v_{2}\right)-\hat{r}\left(v_{1}\right)}{v_{2}-v_{1}}.$
This adapted LIV approach is computationally more convenient than the separate estimation procedure. Furthermore, if we accept a parametric specification on $E\left[U_{1}-U_{0}\left|V=v\right.\right]$ and thus on $q\left(p\right)$, the adapted LIV approach can be implemented through a parametric least squares regression.
In summary, the identification based on functional form can accommodate most of the frequently-used estimation procedures in a typical IV model.
@@ 5 Empirical application
In this section, we revisit the heterogeneous long-term effects of the Head Start program on educational attainment and labor market outcomes by using the proposed MTE method. Head Start, which began in 1965, is one of the largest early child care programs in the United States. The program is targeted at children from low-income families and can provide such children with preschool, health, and nutritional services. Currently, Head Start serves more than a million children, at an annual cost of 10 billion dollars. As a federally funded large-scale program, Head Start has encountered concerns about its effectiveness and thus spawned numerous studies to evaluate its educational and economic effects on the participants. Early studies focused mainly on short-term benefits and found that participation in Head Start is associated with improved test scores and reduced grade repetition at the beginning of primary schooling. However, such benefits seem to fade out during the upper primary grades (Currie, 2001). Garces et al. (2002) provided the first empirical evidence for the longer-term effects of Head Start on high school completion, college attendance, earnings, and crime. Since then, the literature has shifted its focus to the medium- and long-term or intergenerational (e.g., Barr and Gibbs, 2022) gains of Head Start enrollment.
@@
We first verify the credibility of Assumptions S and NL for the Head Start data. Table 1 lists the support of the nonparametrically estimated propensity score for all data cells that are split by different values of discrete covariates. We find that the 34th data cell has a full support on the unit interval. Therefore, Assumption S must hold if we choose the corresponding values (32 years old, female, black race, parental education being college or higher) as the benchmark. We then verify Assumption NL for the 34th data cell. Note that only one continuous covariate exists in the data, that is, family income in 1978. Hence, we invoke Assumption NL1, which requires the propensity score function to have at least one stationary point. The left panel of Figure 1 plots the estimated propensity score as a univariate function of the continuous covariate for the 34th data cell. We find that the propensity score is highly nonlinear and nonmonotonic in family income in 1978, possibly owing to the parents’ various self-selection on participation in the program. So Assumption NL1 is fulfilled.
@@
Furthermore, the largest data cell (i.e., the 29th data cell with 175 observations) can be considered an alternative candidate for the benchmark (32 years old, male, white race, parental education being high school or lower). Nevertheless, closer examination of Assumption S is necessary for this data cell, as the support of the propensity score is now limited to $\left[0,0.194\right]$, which is totally separate from the supports for some other data cells. To this end, we need to find a distinct value of each discrete variable such that the propensity score has overlapping support with $\left[0,0.194\right]$. For example, if the value of age is altered from 32 to 30 (or 31, 33, 34) while the other discrete variables remain unaltered, we will go to the 5th (or 17th, 41st, 53rd) data cell with support $\left[0,0.177\right]$ (or $\left[0,0.242\right],\left[0,0.154\right],\left[0,0.204\right]$), which is overlapped with $\left[0,0.194\right]$ as required. Altering the values of gender, race, and parental education leads to the 35th, 25th (or 27th), and 30th data cells with supports $\left[0,0.526\right]$, $\left[0,0.771\right]$ (or $\left[0,0.623\right]$), and $\left[0,0.061\right]$, respectively, which all meet the overlapping condition as well. Therefore, Assumption S holds for the largest data cell. The validation of Assumption NL1 for this data cell can be demonstrated as well by the existence of stationary points of the estimated propensity score function, as illustrated in the right panel of Figure 1.
@@
We next turn to the estimation. In the first step, we nonparametrically regress the treatment variable (Head Start) on all of the covariates to generate propensity scores for the sample, by employing the kernel estimation method in (10) and (11) with the rule-of-thumb bandwidth. Figure 2 plots the frequency distribution of the estimated propensity score by treatment status. The figure shows that the propensity score in our sample follows a bimodal distribution, with the main peak being at approximately 0.5 for the participants and approximately 0.1 for the nonparticipants. To reduce the potential impact of outliers, we trim the observations of the 1% smallest and 1% largest propensity scores in the following steps. This trimming leads to a common support ranging from 0 to 0.6, as indicated by the two dashed vertical lines. Given the estimated propensity score, we then estimate in sequence the linear coeffecients, MTE, and summary treatment effect measures including ATE, TT, and TUT. We adopt the separate estimation procedure, since it generally exploits more identifying information behind the data than the LIV procedure (Brinch et al., 2017).
@@
Figure 3 shows the estimated MTE under Heckman’s normal specification that $E\left[\left.U_{d}\right|V=v\right]=\rho_{d}\Phi^{-1}\left(v\right)$ for $d=0,1$, where $\Phi\left(\cdot\right)$ is the CDF of normal distribution, which can be derived by assuming that $\left(U_{d},V\right)$ follows a bivariate normal distribution with covariance $\rho_{d}$. The MTE curves, evaluated at the mean values of $X$, relate the unobserved component $U_{1}-U_{0}$ of the treatment effect to the unobserved component $V$ of the treatment choice. A high value of $V$ implies a low probability of treatment; thus, we interpret $V$ as resistance to participation in Head Start. The MTE curve for wage income plotted in panel B decreases with resistance, revealing a pattern of selection on gains, as expected. In other words, based on the unobserved characteristics, the children who were most likely to enroll in Head Start benefitted the most from the program in terms of their labor market outcome. However, when the outcome is educational attainment in panel A, the positive slope of the MTE curve points to a pattern of reverse selection on gains. In consequence, the TUT exceeds the ATE, which in turn exceeds the TT. The same pattern was observed by Cornelissen et al. (2018) when estimating school-readiness return to a preschool program in Germany. This phenomenon may be attributed partly to the fact that parents have their own objectives in deciding childcare arrangements. Nevertheless, the disagreement between selection patterns for the two outcomes raises concern about the possible functional form misspecification of the normal MTE, which is strictly restricted to be monotonic in the resistance to treatment. Therefore, we consider a nonparametric specification for $E\left[\left.U_{d}\right|V=v\right]$ and implement a semiparametric separate estimation procedure.
@@
for potential outcomes $Y_{1}$ and $Y_{0}$, which we plot in Figure 5. Panel A sheds light on the significantly positive effect of Head Start on education for the respondents with medium-level resistance to treatment, revealing that their gains from the program are driven mainly by the remarkably low educational attainment when untreated. Panel B leads to similar results for wage income, where significant gain emerges for low values of $V$. Moreover, the relatively flat curve of potential wage income in the treated state implies that early childcare attendance serves as an equalizer that diminishes the intergroup difference in the labor market outcome.
@@
where the last equality follows from the independence of $X$ and $V$ and from the conditional mean independence of $X$ and $U_{d}$ given $V$. Plugging in proper estimates of each component in the above equations, we obtain the estimated marginal response functions and plot them in Figure 6. In particular, the marginal observed outcome curves relate the individuals with significant positive returns (with medium $V$ in Panel A and small $V$ in Panel B) to those with low education and low wage income, thereby bridging our results on the MTE and those of De Haan and Leuven (2020) on the distributional treatment effect.
6 Conclusion
We propose a novel method for defining, identifying, and estimating the MTE in the absence of IVs. Our MTE model allows all the covariates to be correlated to the structural treatment error. In this model, we define the MTE based on a reduced-form treatment error that (i) is uniformly distributed on the unit interval, (ii) is statistically independent of covariates, and (iii) has several economic meanings. The independence property facilitates the identification of our defined MTE. We provide sufficient conditions under which the MTE can be point identified based on functional form. The conditions are standard in a certain sense. The conditional mean independence assumption is equivalent to the separability assumption commonly imposed in the MTE literature, and is implied by and much weaker than the full independence assumption. The linearity and nonlinearity assumptions are the foundation of identification based on functional form, and can make sense to most empirical studies. We prove the identification by using a construction method. Our identification strategy allows the adaptation of most of the existing estimation procedures for conventional IV-MTE models, such as separate estimation, LIV estimation, parametric estimation, and semiparametric estimation. For the empirical application, we evaluate the MTE of the Head Start program on long-term education and labor market outcomes, in which an IV for Head Start participation is difficult to acquire. We find significant positive effects for individuals with medium-level or low resistance to treatment, and substantial heterogeneity exists.
Appendix A Proof of Theorem 1
Denote $m_{d}\left(x\right)=E\left[Y\left|X=x,D=d\right.\right]$ and $m_{d0}\left(x^{C}\right)=m_{d}\left(x^{C},0\right)$, $d=0,1$. Note that $\pi\left(x\right)$ and $m_{d}\left(x\right)$, and thus $\pi_{0}\left(x^{C}\right)$ and $m_{d0}\left(x^{C}\right)$, are identified functions because they are conditional expectations of observed variables. We first consider identifying $\beta_{d}^{C}$, the coefficients of continuous covariates, from $\pi_{0}\left(x^{C}\right)$ and $m_{d0}\left(x^{C}\right)$. By equation (7), we have
$m_{d0}\left(x^{C}\right)=x^{C\prime}\beta_{d}^{C}+g_{d}\left(\pi_{0}\left(x^{C}\right)\right).$ (A.1)
When $\dim\left(X^{C}\right)=1$, by differentiating both sides of (A.1) at $\tilde{x}^{C}$ and invoking Assumption NL1, we obtain

### v3

@@ 4 Estimation
The kernel-weighted pairwise difference estimator has the advantage of having a closed-form expression, so we need not solve any formidable optimization problems. However, it faces the challenging problem of bandwidth selection as most semiparametric estimation methods. Alternatively, we can consider imposing a parametric specification on the unobservable heterogeneity of the MTE such that $E\left[\left.U_{d}\right|V=v\right]=E\left[\left.U_{d}\right|V=v;\theta_{d}\right]$ for finite dimensional $\theta_{d}$ (Heckman and Vytlacil, 2005), e.g., the polynomial specification such that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}v^{j}$ or the normal polynomial specification such that $E\left[\left.U_{d}\right|V=v\right]=\sum_{j=0}^{J}\theta_{dj}\Phi^{-j}\left(v\right)$, for a fixed $J$. In the latter, setting $J=1$ will match Heckman’s normal sample selection model. Under the parametric restriction, the second step becomes a global regression $E\left[Y\left|X,D=d\right.\right]=X^{\prime}\beta_{d}+g_{d}\left(P;\theta_{d}\right)$ with parameterized selectivity bias correction term $g_{d}\left(p;\theta_{d}\right)$ for each $d$; therefore, the tuning of $h_{2}$ and $h_{3}$ is circumvented. Another advantage of a parametrically specified second step is the lower computational burden relative to the pairwise difference estimator that is defined by double summation which requires a squared amount of calculation (Pan and Xie, 2023). The simulation in Appendix G shows that MTE estimators with parametric second-step perform well when the specification is correct or nearly correct, while the semiparametric MTE estimator performs robustly across different designs.
@@
[…]
@@
$q\left(p\right)=pE\left[U_{1}-U_{0}\left|V\leq p\right.\right]=\int_{0}^{p}E\left[U_{1}-U_{0}\left|V=v\right.\right]dv.$
Note that since
@@
$\left(\begin{array}[]{c}\check{\beta}_{0}\\ \check{\delta}\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}k_{2}\left(\frac{\hat{P}_{i}-\hat{P}_{j}}{h_{2}}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)\left(\begin{array}[]{c}X_{i}-X_{j}\\ \hat{P}_{i}X_{i}-\hat{P}_{j}X_{j}\end{array}\right)^{\prime}\right]^{-1}$
@@
where $\check{\delta}$ is an estimator for $\beta_{1}-\beta_{0}$ according to (15). Given $\check{\beta}_{0}$ and $\check{\delta}$, we apply the local linear method as well, yielding
$\left(\begin{array}[]{c}\check{r}\left(p\right)\\ \check{q}^{\left(1\right)}\left(p\right)\end{array}\right)$ $=$ $\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)^{\prime}\right]^{-1}$
$\cdot\left[\sum_{i=1}^{n}k_{3}\left(\frac{\hat{P}_{i}-p}{h_{3}}\right)\left(\begin{array}[]{c}1\\ \hat{P}_{i}-p\end{array}\right)\left(Y_{i}-X_{i}^{\prime}\check{\beta}_{0}-\hat{P}_{i}X_{i}^{\prime}\check{\delta}\right)\right],$
where $\check{r}\left(p\right)$ is an estimator for $\alpha_{0}+q\left(p\right)$. Then, we construct the instrument-free LIV estimators for the MTE and other treatment parameters as
$\check{\Delta}^{\text{MTE}}\left(x,v\right)$ $=$ $x^{\prime}\check{\delta}+\check{q}^{\left(1\right)}\left(v\right),$
$\check{\Delta}^{\text{ATE}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\check{r}\left(1\right)-\check{r}\left(0\right),$
$\check{\Delta}^{\text{TT}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(\hat{\pi}\left(x\right)\right)-\check{r}\left(0\right)}{\hat{\pi}\left(x\right)},$
$\check{\Delta}^{\text{TUT}}\left(x\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(1\right)-\check{r}\left(\hat{\pi}\left(x\right)\right)}{1-\hat{\pi}\left(x\right)},$
$\check{\Delta}^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $x^{\prime}\check{\delta}+\frac{\check{r}\left(v_{2}\right)-\check{r}\left(v_{1}\right)}{v_{2}-v_{1}}.$
[…]
In summary, the identification based on functional forms can accommodate most of the frequently-used estimation procedures in IV-based MTE models. However, since it is infeasible to manipulate the nonlinear variation, our identification result does not allow for the counterfactual analysis via estimating the policy-relevant treatment effect (PRTE) which is also an important application of IV-based MTE models (Heckman and Vytlacil, 2005).
@@ 5 Empirical application
In this section, we revisit the heterogeneous long-term effects of the Head Start program on educational attainment and labor market outcomes by using the proposed IV-free MTE method. Head Start, which began in 1965, is one of the largest early child care programs in the United States. The program is targeted at children from low-income families and can provide such children with preschool, health, and nutritional services. Currently, Head Start serves more than a million children, at an annual cost of 10 billion dollars. As a federally funded large-scale program, Head Start has encountered concerns about its effectiveness and thus spawned numerous studies to evaluate its educational and economic effects on the participants. Early studies focused mainly on short-term benefits and found that participation in Head Start is associated with improved test scores and reduced grade repetition at the beginning of primary schooling. However, such benefits seem to fade out during the upper primary grades (Currie, 2001). Garces et al. (2002) provided the first empirical evidence for the longer-term effects of Head Start on high school completion, college attendance, earnings, and crime. Since then, the literature has shifted its focus to the medium- and long-term or intergenerational (e.g., Barr and Gibbs, 2022) gains of Head Start enrollment.
@@
We first verify the credibility of Assumptions S and NL for the Head Start data. Table F.2 in Appendix F lists the support of the nonparametrically estimated propensity score for all data cells that are split by different values of discrete covariates. We find that the 34th data cell has a full support on the unit interval. Therefore, Assumption S must hold if we choose the corresponding values (32 years old, female, black race, parental education being college or higher) as the benchmark. We then verify Assumption NL for the 34th data cell. Note that only one continuous covariate exists in the data, that is, family income in 1978. Hence, we invoke Assumption NL1, which requires the propensity score function to have at least one stationary point. The left panel of Figure 1 plots the estimated propensity score as a univariate function of the continuous covariate for the 34th data cell. We find that the propensity score is highly nonlinear and nonmonotonic in family income in 1978, possibly owing to the parents’ various self-selection on participation in the program. So Assumption NL1 is fulfilled.
@@
According to our identification result, it is enough to find one benchmark value of $X^{D}$ for which the support of the propensity score satisfies Assumption S and the propensity score function satisfies Assumption NL, as fulfilled above. In another application, however, the chances are that the propensity score function corresponding to the fully supported data cell violates Assumption NL or even there is no fully supported data cell. If this is the case, we need to try values of $X^{D}$ for which the verification of Assumption S may be not so trivial. As an example, we consider an additional data cell and retest Assumptions S and NL. We choose the largest data cell (i.e., the 29th data cell with 175 observations) and reset the benchmark to be 32 years old, male, white race, and high school or lower parental education. Now closer examination of Assumption S is necessary, as the support for this data cell is limited to $\left[0,0.194\right]$, which is totally separate from the supports for some other data cells. To this end, we need to find a distinct value of each discrete variable such that the propensity score has overlapping support with $\left[0,0.194\right]$. For example, if the value of age is altered from 32 to 30 (or 31, 33, 34) while the other discrete variables remain unaltered, we will go to the 5th (or 17th, 41st, 53rd) data cell with support $\left[0,0.177\right]$ (or $\left[0,0.242\right],\left[0,0.154\right],\left[0,0.204\right]$), which is overlapped with $\left[0,0.194\right]$ as required. Altering the values of gender, race, and parental education leads to the 35th, 25th (or 27th), and 30th data cells with supports $\left[0,0.526\right]$, $\left[0,0.771\right]$ (or $\left[0,0.623\right]$), and $\left[0,0.061\right]$, respectively, which all meet the overlapping condition as well. Therefore, Assumption S holds for the largest data cell. The validation of Assumption NL1 for this data cell can be demonstrated as well by the existence of stationary points of the estimated propensity score function, as illustrated in the right panel of Figure 1.
@@
We next turn to the estimation. In the first step, we nonparametrically regress the treatment variable (Head Start) on all of the covariates to generate propensity scores for the sample, by employing the kernel estimation method in (12) with the Gaussian kernel and the rule-of-thumb bandwidth as given in Table F.1. Figure 2 plots the frequency distribution of the estimated propensity score by treatment status. The figure shows that the propensity score in our sample follows a bimodal distribution, with the main peak being at approximately 0.5 for the participants and approximately 0.1 for the nonparticipants. To reduce the potential impact of outliers, we trim the observations of the 1% smallest and 1% largest propensity scores in the following steps. This trimming leads to a common support ranging from 0 to 0.6, as indicated by the two dashed vertical lines. Given the estimated propensity score, we then estimate in sequence the linear coefficients $\beta_{1}$ and $\beta_{0}$, MTE, and summary treatment effect measures including ATE, TT, and TUT. Since the separate estimation procedure exploits more identifying information behind the data than the LIV procedure in general (Brinch et al., 2017), we focus on the results of the separate estimation.
@@
Figure 3 shows the estimated MTE under Heckman’s normal specification in which $E\left[\left.U_{d}\right|V=v\right]=\rho_{d0}+\rho_{dV}\Phi^{-1}\left(v\right)$ for $d=0,1$, where $\Phi\left(\cdot\right)$ is the CDF of standard normal distribution, which can be derived by assuming that $\left(U_{d},U\right)$ follows a bivariate normal distribution as in Example 3. The MTE curves, evaluated at the mean values of $X$, relate the unobserved component $U_{1}-U_{0}$ of the treatment effect to the unobserved component $V$ of the treatment choice. A high value of $V$ implies a low probability of treatment; thus, we interpret $V$ as resistance to participation in Head Start. The MTE curve for wage income plotted in panel B decreases with resistance, revealing a pattern of selection on gains, as expected. In other words, based on the unobserved characteristics, the children who were most likely to enroll in Head Start benefitted the most from the program in terms of their labor market outcome. However, when the outcome is educational attainment in panel A, the positive slope of the MTE curve points to a pattern of reverse selection on gains. In consequence, the TUT exceeds the ATE, which in turn exceeds the TT. The same pattern was observed by Cornelissen et al. (2018) when estimating school-readiness return to a preschool program in Germany. This phenomenon may be attributed partly to the fact that parents have their own objectives in deciding childcare arrangements. Nevertheless, the disagreement between selection patterns for the two outcomes raises concern about the possible functional form misspecification of the normal MTE, which is strictly restricted to be monotonic in the resistance to treatment. Therefore, we consider a nonparametric specification for $E\left[\left.U_{d}\right|V=v\right]$ and implement a semiparametric separate estimation procedure.
@@
6 Conclusion
We propose a novel method for defining, identifying, and estimating the MTE in the absence of IVs. Our MTE model allows all the covariates to be correlated to the structural treatment error. In this model, we define the MTE based on a reduced-form treatment error that (i) is uniformly distributed on the unit interval, (ii) is statistically independent of covariates, and (iii) has several economic meanings. The independence property facilitates the identification of our defined MTE. We provide sufficient conditions under which the proposed IV-free MTE can be point identified based on functional forms. The conditions are standard in a certain sense. The conditional mean independence assumption is equivalent to the separability assumption commonly imposed in the MTE literature, and is implied by and much weaker than the full independence assumption. The linearity and nonlinearity assumptions are the foundation of identification based on functional forms, and can make sense in most empirical studies. We prove the identification by using a construction method. Our identification strategy allows the adaptation of most of the existing estimation procedures for the IV-based MTE, such as separate estimation, LIV estimation, parametric estimation, and semiparametric estimation. For the empirical application, we evaluate the MTE of the Head Start program on long-term education and labor market outcomes, in which an IV for Head Start participation is difficult to acquire. We find significant positive effects for individuals with medium-level or low resistance to treatment, and substantial heterogeneity exists.
Acknowledgements
We wish to thank the Editor and two anonymous reviewers for their valuable comments and suggestions. Financial supports for this research are provided through the National Social Science Foundation of China (Grant 24FTJB010) and the National Natural Science Foundation of China (Grants 72173142, 72034006, 72342034, 72173083).
Appendix A Additional results of the empirical application
@@
for potential outcomes $Y_{1}$ and $Y_{0}$, which we plot in Figure A.1. Panel A sheds light on the significantly positive effect of Head Start on education for the respondents with medium-level resistance to treatment, revealing that their gains from the program are driven mainly by the remarkably low educational attainment when untreated. Panel B leads to similar results for wage income, where significant gain emerges for low values of $V$. Moreover, the relatively flat curve of potential wage income in the treated state implies that early childcare attendance serves as an equalizer that diminishes the intergroup difference in the labor market outcome.
Figure A.1: Semiparametric estimates of marginal structural functions
Notes: The marginal structural function refers to the expected potential outcome conditional on the unobserved resistance to treatment. Panels A and B depict the estimated marginal structural function curves for education outcome and labor market outcome, respectively. The 90% confidence interval is based on bootstrapping with 1,000 replications.
@@
where the last equality follows from the independence of $X$ and $V$ and from the conditional mean independence of $X$ and $U_{d}$ given $V$. Plugging in proper estimates of each component in the above equations, we obtain the estimated marginal response functions and plot them in Figure A.2. In particular, the marginal observed outcome curves relate the individuals with significant positive returns (with medium $V$ in Panel A and small $V$ in Panel B) to those with low education and low wage income, thereby bridging our results on the MTE and those of De Haan and Leuven (2020) on the distributional treatment effect.
Figure A.2: Semiparametric estimates of marginal response functions
Notes: This figure plots the (conditional) expectation of the observed outcome (solid line on the left y-axis), and the conditional probability of participation (dashed line on the right y-axis). Panels A and B illustrate the estimated marginal response function curves for education outcome and labor market outcome, respectively. The 90% confidence interval is based on bootstrapping with 1,000 replications.
Appendix B Proof of Theorem 1
[…]
$m_{d0}\left(x^{C}\right)=x^{C\prime}\beta_{d}^{C}+g_{d}\left(\pi_{0}\left(x^{C}\right)\right).$ (B.1)
When $\dim\left(X^{C}\right)=1$, by differentiating both sides of (B.1) at $\tilde{x}^{C}$ and invoking Assumption NL1, we obtain